OpenStudy (anonymous):

Solve the system of equations 4x+5y=-6 4x-7y=18

OpenStudy (anonymous):

subtract two systems and omit x then you have 12y=-24 so y is found and then looking for x

OpenStudy (anonymous):

so is it y=-2 and x=1

OpenStudy (lgbasallote):

seems right to me

OpenStudy (anonymous):

4x - 4x 5y - (-7y) = -6 -18 12 y = -24 y = -24/12 y = -2 sub y=-2 back into one of the previous equations lets use 4x + 5y = -6 4x + 5(-2) = -6 4x -10 = -6 4x = -6 + 10 4x = 4 x = 4/4 x = 1 yes u are right!

OpenStudy (anonymous):

yay,i cn only solve these thingys with a help of how 2 start 2 solve the equation but i shld no it bc im in 10th gnna be in 11th after i take a test tomm so i shld problley learn to solve it

OpenStudy (anonymous):

^^^^^^You should probably learn how to spell to......hahah jkjk